# Stick-breaking Attention

### v1

@@ 5.1 Multi-Query Repeated Associative Recall Task
While stick-breaking can also solve the MQAR task, we formulated a different version of this toy task and tested it on both softmax attention and stick-breaking. As before, each instance of the task has an initial assignment of values to variables. However, in the query sequence, the same variable can be queried multiple times, and each query is then followed by a variable assignment, which successive queries must recall. We refer to this task as multi-query repeated associative recall (MQRAR).
@@ 5.3 Model Pretraining
We pretrain the 1B and 3B models in this section using a two-stage training scheme in Hu et al. (2024) and the Power learning rate schedule (Shen et al., 2024). In the first stage, there is a warmup for the learning rate to 0.01, then we apply Power decay. Our training corpus has 1T tokens and mixes large-scale open-source datasets of medium quality with permissive licenses. In the second stage, we exponentially decay the learning rate to zero. The stage 2 training corpus is a mix of stage 1 data and a small amount of high-quality open-source and synthetic corpora with permissive licenses. The training batch size is 1024 and uses padding-free sequence packing for training in Dolomite Engine (Mishra, 2024). We evaluate the pretrained models on language model tasks and multiple-choice tasks from LM evaluation Harness (Gao et al., 2023). The multiple-choice tasks include: grade-school science questions (ARC; Clark et al. 2018), common sense reasoning (Hellaswag; Zellers et al. 2019), open book question answering (OpenBookQA; Mihaylov et al. 2018), physical questions (PIQA; Bisk et al. 2020), reading comprehension (RACE; Lai et al. 2017), and Winograd schema task (Winogrande; Sakaguchi et al. 2021). Table 2 shows the performance.
Overall, we outperform our own pretrained standard attention models that are trained on the same settings. We perform better on average, and attain better perplexity on Wikitext. On the 1B experiment, we find that the stick-breaking without remainder performs better, so we only pretrained a 3B parameter model on this setting. Subsequent experiments are also run only on the non-remainder setting. We also evaluated the models on MMLU with 0-shot and 5-shot settings. Due to the inductive bias of stick-breaking, we believed that stick-breaking would perform better on MMLU in a few-shot setting as it would not be distracted by the few-shot examples provided in the context.
@@
We have included the the Qwen1.5-4B and Gemma2-2B models for comparison, and our 3B model underperforms Gemma2-2B, a smaller model, while it outperforms Qwen1.5-4B. This may be attributable to a variety of factors, including data mixture or training hyperparameters. Finally, we evaluate our 3B model on the GSM8K dataset (Cobbe et al., 2021). Interestingly, the 5-shot setting underperforms standard attention while CoT with 8-shot sees a 5.5% improvement.
@@
Figure 6: Results on the RULER benchmark. Without further finetuning for long context, we evaluate the 1B models on the RULER tasks. Subfigures: 6(a) shows the average over all 13 task, 6(b) is the average across the various NIAH tasks, 6(c) is the result for the variable tracking task.
@@ 6 Conclusion & Future Work
The drawbacks in efficiency can be improved in future work by making similar optimisations that Flash Attention (Dao et al., 2022; Dao, 2023) made for speedups. These include making full use of features in CUDA, and by cache retrieval optimisations. We believe there is plenty of room for improvement in computation efficiency that can be made in future versions of stick-breaking. Overall, we find stick-breaking attention to be a promising replacement for Softmax + RoPE in Transformer models.
@@ Acknowledgments
Appendix A 1B Model Ruler Results

### v2

@@ 5.1 Multi-Query Repeated Associative Recall Task
While stick-breaking can also solve the MQAR task, we formulated a different version of this toy task and tested it on both softmax attention and stick-breaking. As before, each instance of the task has an initial assignment of values to variables. However, in the query sequence, the same variable can be queried multiple times, and each query is then followed by a variable assignment, which successive queries must recall. We refer to this task as multi-query repeated associative recall (MQRAR). This setting can be analogous to some scenarios in programming where variables are repeatedly updated, and it where it is useful to understand the current state of the variable assignment. The following is an example sequence of the task:
@@ 5.3 Language Model Pretraining
We pretrain the 1B and 3B models in this section using a two-stage training scheme in Hu et al. (2024) and the Power learning rate schedule (Shen et al., 2024). In the first stage, there is a warmup for the learning rate to 0.01, then we apply Power decay. Our training corpus has 1T tokens and mixes large-scale open-source datasets of medium quality with permissive licenses. In the second stage, we exponentially decay the learning rate to zero. The stage 2 training corpus is a mix of stage 1 data and a small amount of high-quality open-source and synthetic corpora with permissive licenses. The training batch size is 1024 and uses padding-free sequence packing for training in Dolomite Engine (Mishra, 2024). We evaluate the pretrained models on language model tasks and multiple-choice tasks from LM evaluation Harness (Gao et al., 2023). The multiple-choice tasks include: grade-school science questions (ARC; Clark et al. 2018), common sense reasoning (Hellaswag; Zellers et al. 2019), open book question answering (OpenBookQA; Mihaylov et al. 2018), physical questions (PIQA; Bisk et al. 2020), reading comprehension (RACE; Lai et al. 2017), and Winograd schema task (Winogrande; Sakaguchi et al. 2021). Table 2 shows the performance. Overall, we outperform our own pretrained standard attention models that are trained on the same settings. We perform better on average, and attain better perplexity on Wikitext.
@@
We also evaluated the models on MMLU with 0-shot and 5-shot settings. For the 1B models, we have included the results for TinyLlama (Zhang et al., 2024) Due to the inductive bias of stick-breaking, we believed that stick-breaking would perform better on MMLU in a few-shot setting as it would not be distracted by the few-shot examples provided in the context.
We have included the Qwen1.5-4B and Gemma2-2B models for comparison, and our 3B model underperforms Gemma2-2B, a smaller model, while it outperforms Qwen1.5-4B.
Finally, we evaluate our 3B model on the GSM8K dataset (Cobbe et al., 2021). Interestingly, the 5-shot setting underperforms standard attention while CoT with 8-shot sees an improvement of 5.5%. The improvements in MMLU and GSM8K on a few-shot setting may suggest stick-breaking has a better inductive bias for in-context learning, particularly in the formats that these evaluation benchmarks use.
@@
Figure 6: Results on the RULER benchmark. Without further finetuning for long context, we evaluate the 1B models on the RULER tasks. Subfigures: 6(a) shows the average over all 13 tasks, 6(b) is the average across the various NIAH tasks, 6(c) is the result for the variable tracking task.
@@ 6 Conclusion & Future Work
The drawbacks in efficiency can be improved in future work by making similar optimisations that Flash Attention (Dao et al., 2022; Dao, 2023) made for speedups. These include making full use of features in CUDA, and by cache retrieval optimisations. We believe there is plenty of room for improvement in computation efficiency that can be made in future versions of stick-breaking. We also think there can be stick-breaking-specific modifications to Transformers that can augment the Stick-breaking Transformer, and we have initial results in Appendix C. Overall, we find stick-breaking attention to be a promising replacement for Softmax + RoPE in Transformer models.
@@ Acknowledgments
Appendix A Derivation of Log-space Formulation
$\sigma(x)$ is the sigmoid function,
$\sigma(x)$ $=\frac{1}{1+\exp(-x)}=\frac{\exp(x)}{1+\exp(x)}$ (7)
$1-\sigma(x)$ $=\frac{1}{1+\exp(x)}$ (8)
Since $\beta_{i,j}=\sigma(z_{i,j})$,
$\log\beta_{i,j}$ $=z_{i,j}-\log\left(1+\exp(z_{i,j})\right)$ (9)
$\log\left(1-\beta_{i,j}\right)$ $=\log\frac{1}{1+\exp(z_{i,j})}=-\log\left(1+\exp(z_{i,j})\right)$ (10)
We can substitute these back in,
${\bm{A}}_{i,j}$ $=\exp\left(\log\beta_{i,j}+\sum_{k=i+1}^{j-1}\log\left(1-\beta_{k,j}\right)\right)$ (11)
$=\exp\left(z_{i,j}-\log\left(1+\exp(z_{i,j})\right)-\sum_{k=i+1}^{j-1}\log\left(1+\exp(z_{k,j})\right)\right)$ (12)
$=\exp\left(z_{i,j}-\sum_{k=i}^{j-1}\log\left(1+\exp(z_{k,j})\right)\right)$ (13)
Appendix B 1B Model RULER Results
@@ Appendix A Derivation of Log-space Formulation
Appendix C Stick-breaking Specific Transformer Modifications
While we demonstrate the effectiveness of stick-breaking as a drop-in replacement for softmax attention in the main paper, found several minor modifications to the standard Transformer architecture can help enhance the performance of the Stick-breaking Transformer on various tasks.
Table 5: In these experiments, we use $n_{\mathrm{layer}}=40$, uses Grouped Query Attention with $n_{\mathrm{head}}=12$ and 4 key and value heads, and $d_{\mathrm{head}}=128$. For the 6.5B model, we use $n_{\mathrm{embd}}=1536$, with 64 experts and $k=8$, with each expert having a hidden size of 512. For the 28.6B model, we use $n_{\mathrm{embd}}=4096$, with 72 experts and $k=12$, with each expert having a hidden size of 768.
RB GN ARC Hella. OBQA PIQA Wino. MMLU GSM8K MATH Avg.
Accuracy (normalised) Accuracy 8-shot, cot 4-shot
1.2B parameters, 1.25T tokens
Softmax 55.8 66.7 38.4 76.0 63.9 28.3 27.0 7.6 45.5
SB 54.0 66.9 39.6 76.8 65.5 37.3 29.6 8.2 47.2
54.1 66.6 39.8 76.2 66.9 39.6 29.2 7.8 47.5
56.4 66.7 38.4 76.7 64.8 39.8 29.3 8.2 47.5
1B activation / 6.5B parameters, 2T tokens
Softmax 66.3 76.9 44.6 80.7 72.5 55.5 50.8 21.6 58.6
SB 68.7 77.9 46.6 80.5 74.4 56.9 55.3 22.5 60.4
6B activation / 28.6B parameters, 3T tokens
Softmax 76.8 84.4 51.2 84.1 80.1 67.7 70.4 34.5 68.6
SB 77.2 84.7 51.4 83.7 81.5 71.2 71.9 35.6 69.6
Remainder Bias (RB)
As $\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\leq 1$, we experimented with using remaining attention weight is assigned to an embedding ${\bm{r}}$, by modifying the output of the attention layer as follows:
${\bm{o}}_{j}$ $=\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\cdot{\bm{v}}_{i}+\left(1-\sum_{i=1}^{j-1}{\bm{A}}_{i,j}\right)\cdot{\bm{r}}$ (14)
Each head has its corresponding ${\bm{r}}$ embedding, resulting in a matrix of additional parameters of $n_{\mathrm{head}}\cdot d_{\mathrm{head}}$. We refer to this as remainder bias. This is similar to an attention sink token (Xiao et al., 2023). The authors showed that having an attention sink token of $\mathbf{0}$ resulted in poorer performance than the standard attention, and saw better performance when a learnable attention sink embedding was used. In our experiments, we saw improvements when the remainder bias was used.
Group Norm (GN)
Due to the unnormalised weight of stick-breaking attention, the resulting ${\bm{o}}_{j}$ can have varying norm magnitudes. We experimented with applying a head-wise norm, which was implemented with Group Norm. This was also done with Differential Transformers (Ye et al., 2024), for similar reasons. We found improvements when Group Norm was used in conjunction with RB.
In Table 5, we perform experiments with RB and GN on 1B models where we found improvements on 1B dense models. We then trained a 6.5B and 28.6B parameter MoE model with both GN and RB, and show that they outperform a corresponding Transformer with softmax attention.
C.1 Context window extension
All models are pre-trained with a context window of 4096. To extend their context lengths, we continue training for 6250 steps at a context length of 16k. Our learning rate is scheduled to increase from $10^{-5}$ to $0.000125$ in 150 steps, and then decays exponentially to 0 until 6250 steps. The effective batch size has 4M tokens.
For Softmax+RoPE, we set the dynamic RoPE scaling factor to 4. In Figure , we show that by applying rope scaling, the model can somewhat extrapolate to 16k context lengths, but the log-likelihood spikes as we extend it further.
Applying this method of context window extension to both models, we find that the Softmax+RoPE model improves slightly, while still showing issues with longer context lengths, albeit mitigated slightly. Stick-breaking however exhibits lower NLL as we extend the context window further to 32k and 64k.
The behaviour of RoPE out-of-the-box for length extrapolation is also demonstrated in Section 5.2. We also show that training at longer sequences does not mitigate the out-of-distribution performance of Softmax+RoPE models. Here, we show that while naive continued fine-tuning helps, the improvement on long-context windows is not as substantial as with stick-breaking. This suggests better context window extension behaviour and further length extrapolation for the stick-breaking attention mechanism.
Figure 8: Stick-breaking and Softmax+RoPE context window extrapolation performance on longer context lengths in the Pile 10k evaluation dataset.